# Mineral Detection of Cosmic-Ray Boosted Dark Matter

### v1

@@ Abstract
We present the first dedicated analysis of cosmic-ray dark matter (CRDM) in paleo detectors. Owing to their large kinetic energies, CRDM particles generate nuclear-recoil tracks that extend to substantially larger lengths than those produced by dominant backgrounds from neutrinos and intrinsic radioactivity. Combined with the ultra-large effective geological exposure of $\mathcal{O}(10^{5})~\mathrm{t\,yr}$, paleo detectors provide a uniquely sensitive probe of sub-GeV DM. Considering both constant and vector-mediator interactions, we find that paleo detectors improve the sensitivity to the DM–proton scattering cross section by one to two orders of magnitude compared with the latest XENONnT limits.
@@
To overcome this limitation, several scenarios involving “boosted” DM populations have been proposed Agashe et al. (2014); Kouvaris (2015); An et al. (2018); Emken et al. (2018); Bringmann and Pospelov (2019); Wang et al. (2022); Granelli et al. (2022). A particularly robust and unavoidable mechanism was identified in Ref. Bringmann and Pospelov (2019), where a small fraction of the Galactic DM halo is accelerated to relativistic energies through elastic scattering with high-energy cosmic rays (CRs). This component, commonly referred to as cosmic-ray boosted DM (CRDM), acquires sufficient kinetic energy to generate observable signals in terrestrial detectors, thereby enabling sensitivity to DM–nucleon interactions well below the GeV mass scale.
@@
CRDM flux.— For the interaction between DM and nucleons, we consider two benchmark scenarios. The first is a phenomenological model with a constant DM–proton scattering cross section, denoted by $\sigma_{\chi p}$. The second is a simplified model with a massive vector mediator, described by the Lagrangian Guo et al. (2020); Su et al. (2024); Wang (2025)
@@
where $A_{i}$ is the mass number of CR species $i$, $\mu_{\chi i}$ and $\mu_{\chi p}$ are the reduced masses of the DM–nucleus and DM–proton systems, respectively, and $G_{i}(q^{2})$ denotes the nuclear form factor with momentum transfer $q^{2}$. For hydrogen and helium, we adopt the dipole form factor Angeli (2004); Bringmann and Pospelov (2019), while for heavier nuclei we use the Helm form factor Helm (1956); Lewin and Smith (1996); Xia et al. (2022); Wang (2025).
For the vector-mediator scenario, both elastic and deep-inelastic scattering (DIS) processes contribute. The elastic DM–proton scattering cross section is given by Guo et al. (2020); Su et al. (2024); Wang (2025)
$\frac{d\sigma^{\rm el}_{\chi i}}{dT_{\chi}}=2m_{\chi}\frac{A^{2}_{i}g_{p}^{2}g_{\chi}^{2}}{4\pi\beta^{2}}\frac{G_{p}^{2}(Q^{2})}{(Q^{2}+m_{V}^{2})^{2}}\left[1-\frac{\beta^{2}Q^{2}}{Q^{2}_{\rm max}}+\frac{Q^{4}}{8m_{\chi}^{2}E_{i}^{2}}\right],$ (6)
where $g_{p}=3g_{q}$ is the effective DM–proton coupling, $Q^{2}\equiv-q^{2}$ is the positive four-momentum transfer squared, $\beta=|{\bf p}_{i}|/E_{i}$ is the CR’s velocity. The maximal momentum transfer is
@@
where $d\sigma_{\chi j}/dT_{j}$ and the minimal CRDM energy $T_{\chi}^{\rm min}(T_{j})$ follow from Eqs. (6) and (3), with the replacements $\chi\to j$ and $i\to\chi$.
@@
which accounts for the finite readout resolution $\sigma_{x_{T}}$. We consider two representative minerals, Gypsum $\mathrm{Ca(SO_{4})\!\cdot\!2(H_{2}O)}$] and Olivine $\mathrm{Mg}_{1.6}\mathrm{Fe}_{0.4}(\mathrm{SiO}_{4})$], and adopt a benchmark exposure of $100~\mathrm{g\,Gyr}$ Baum et al. (2020a); Drukier et al. (2019); Edwards et al. (2019); Baum et al. (2021b, a); Bramante et al. (2022); Baum and others (2023); Fung et al. (2025). Other exposures can be obtained by simple rescaling.
Only nuclear recoils constitute relevant backgrounds for paleo detectors, since electron recoils do not leave stable tracks. The dominant background sources are neutrinos (solar, atmospheric, and supernova) O’Hare (2020); Baum et al. (2020b), radiogenic neutrons from the 238U decay chain, and recoils induced by 234Th Collar (1996); Snowden-Ifft et al. (1996). For the fiducial $^{238}\mathrm{U}$ concentrations, we adopt the same benchmark values as in Refs. Baum et al. (2020a); Drukier et al. (2019); Edwards et al. (2019); Baum et al. (2021b), namely $10^{-11}$ for Gypsum and $10^{-10}$ for Olivine. Additional backgrounds, such as intrinsic crystal defects, are expected to be distinguishable from recoil tracks Demouchy (2021); Baum et al. (2020b). Signal and background spectra are computed using paleoSpec Baum et al. (2020a); Drukier et al. (2019); Baum et al. (2020b, 2021a).
Figure 2: Binned track length distributions in Gypsum $\mathrm{Ca(SO_{4})\!\cdot\!2(H_{2}O)}$] for CRDM in the vector-mediator model. The red and blue curves show CRDM signals for $m_{\chi}=10^{-5}~\mathrm{GeV}$ and $10^{-3}~\mathrm{GeV}$, respectively, while the black, orange, and green curves denote the dominant nuclear recoil backgrounds from neutrinos, radiogenic neutrons, and $^{238}\mathrm{U}\!\rightarrow\!^{234}\mathrm{Th}+\alpha$ decays. The input parameters are $g_{\chi}=g_{q}=0.1$ and $m_{V}=1~\mathrm{GeV}$.
Fig. 2 shows the binned track length spectrum in Gypsum for the vector-mediator model with $g_{\chi}=g_{q}=0.1$ and $m_{V}=1~\mathrm{GeV}$. The red and blue curves correspond to $m_{\chi}=10^{-5}~\mathrm{GeV}$ and $10^{-3}~\mathrm{GeV}$, respectively. A salient feature is the pronounced excess of CRDM-induced events at large track lengths, where all relevant backgrounds rapidly fall off. This long track regime, which is inaccessible to standard halo DM, provides an efficient and robust handle for background discrimination.

### v2

@@ Abstract
We present the first dedicated analysis of cosmic-ray boosted dark matter (CRDM) in paleo detectors. Owing to their large kinetic energies, CRDM particles generate nuclear-recoil tracks that extend to substantially larger lengths than those produced by dominant backgrounds from neutrinos and intrinsic radioactivity. Combined with the ultra-large effective geological exposure of $\mathcal{O}(10^{5})~\mathrm{t\,yr}$, paleo detectors provide a uniquely sensitive probe of sub-GeV DM. Considering both constant and vector-mediator interactions, we find that paleo detectors improve the sensitivity to the DM–proton scattering cross section by one to two orders of magnitude compared with the latest XENONnT limits.
@@
To overcome this limitation, several scenarios involving “boosted” DM populations have been proposed Agashe et al. (2014); Kouvaris (2015); An et al. (2018); Emken et al. (2018); Bringmann and Pospelov (2019); Wang et al. (2022); Granelli et al. (2022). A particularly robust and unavoidable mechanism was identified in Ref. Bringmann and Pospelov (2019), where a small fraction of the Galactic DM halo is accelerated to relativistic energies through elastic scattering with high-energy cosmic rays (CRs). This component, commonly referred to as cosmic-ray boosted DM (CRDM), acquires sufficient kinetic energy to generate observable signals in terrestrial detectors, thereby enabling sensitivity to DM–proton interactions well below the GeV mass scale.
@@
CRDM flux.— For the interaction between DM and protons, we consider two benchmark scenarios. The first is a phenomenological model with a constant DM–proton scattering cross section, denoted by $\sigma_{\chi p}$. The second is a simplified model with a massive vector mediator, described by the Lagrangian Guo et al. (2020); Su et al. (2024); Wang (2025)
@@
where $A_{i}$ is the mass number of CR species $i$, $\mu_{\chi i}$ and $\mu_{\chi p}$ are the reduced masses of the DM–nucleus and DM–proton systems, respectively, and $G_{i}(Q^{2})$ denotes the nuclear form factor with positive four-momentum transfer squared $Q^{2}$. For hydrogen and helium, we adopt the dipole form factor Angeli (2004); Bringmann and Pospelov (2019), while for heavier nuclei we use the Helm form factor Helm (1956); Lewin and Smith (1996); Xia et al. (2022); Wang (2025).
For the vector-mediator scenario, both elastic and deep-inelastic scattering (DIS) processes contribute. The elastic DM–nucleus $i$ scattering cross section is given by Guo et al. (2020); Su et al. (2024); Wang (2025)
$\frac{d\sigma^{\rm el}_{\chi i}}{dT_{\chi}}=2m_{\chi}\frac{A^{2}_{i}g_{p}^{2}g_{\chi}^{2}}{4\pi\beta^{2}}\frac{G_{i}^{2}(Q^{2})}{(Q^{2}+m_{V}^{2})^{2}}\left[1-\frac{\beta^{2}Q^{2}}{Q^{2}_{\rm max}}+\frac{Q^{4}}{8m_{\chi}^{2}E_{i}^{2}}\right],$ (6)
where $g_{p}=3g_{q}$ is the effective DM–proton coupling, $\beta=|{\bf p}_{i}|/E_{i}$ is the CR’s velocity. The maximal momentum transfer is
@@
[…]
@@
which accounts for the finite readout resolution $\sigma_{x_{T}}$. We consider two representative minerals, Gypsum $\mathrm{Ca(SO_{4})\!\cdot\!2(H_{2}O)}$] and Olivine $\mathrm{Mg}_{1.6}\mathrm{Fe}^{2+}_{0.4}(\mathrm{SiO}_{4})$], and adopt a benchmark exposure of $100~\mathrm{g\,Gyr}$ Baum et al. (2020a); Drukier et al. (2019); Edwards et al. (2019); Baum et al. (2021b, a); Bramante et al. (2022); Baum and others (2023); Fung et al. (2025). Other exposures can be obtained by simple rescaling.
Only nuclear recoils constitute relevant backgrounds for paleo detectors, since electron recoils do not leave stable tracks. The dominant background sources are neutrinos (solar, atmospheric, and supernova) O’Hare (2020); Baum et al. (2020b), radiogenic neutrons from the spontaneous fission or decay chains of 238U, and recoils induced by 234Th Collar (1996); Snowden-Ifft et al. (1996). For the fiducial $^{238}\mathrm{U}$ concentrations, we adopt the same benchmark values as in Refs. Baum et al. (2020a); Drukier et al. (2019); Edwards et al. (2019); Baum et al. (2021b), namely $10^{-11}$ for Gypsum and $10^{-10}$ for Olivine. Additional backgrounds, such as intrinsic crystal defects, are expected to be distinguishable from recoil tracks Demouchy (2021); Baum et al. (2020b). Signal and background spectra are computed using paleoSpec Baum et al. (2020a); Drukier et al. (2019); Baum et al. (2020b, 2021a).
Figure 2: Binned track length distributions in Gypsum $\mathrm{Ca(SO_{4})\!\cdot\!2(H_{2}O)}$] for CRDM in the vector-mediator model. The red and blue curves show CRDM signals for $m_{\chi}=10^{-5}~\mathrm{GeV}$ and $10^{-3}~\mathrm{GeV}$, respectively, while the black, orange, and green curves denote the dominant nuclear recoil backgrounds from neutrinos, radiogenic neutrons, and $^{238}\mathrm{U}\!\rightarrow\!^{234}\mathrm{Th}+\alpha$ decays. The input parameters are $g_{\chi}=g_{q}=0.7$ and $m_{V}=1~\mathrm{GeV}$.
Fig. 2 shows the binned track length spectrum in Gypsum for the vector-mediator model with $g_{\chi}=g_{q}=0.7$ and $m_{V}=1~\mathrm{GeV}$. The red and blue curves correspond to $m_{\chi}=10^{-5}~\mathrm{GeV}$ and $m_{\chi}=10^{-3}~\mathrm{GeV}$, respectively. A salient feature is the pronounced excess of CRDM-induced events at large track lengths, where all relevant backgrounds rapidly fall off. This long track regime, which is inaccessible to standard halo DM, provides an efficient and robust handle for background discrimination.